Hello, please help me find the x in
the roots of x^3+6x^2-4x-24 = 0
thanks!

Answer by ewatrrr(24785)

 
Hi,
x^3+6x^2-4x-24 = 0 | x= 2 is a root, 8 + 24 -8 - 24 = 0
Synthetic Division...
2 1 6 -4 -24
2 16 24
1 8 12 0
Yields:
x^2 +8x + 12 = 0
(x + 6)(x + 2) = 0
x = -6 0r x = -2
Roots: -6, -2, 2